Status Report

ISS On-Orbit Status 29 May 2002

By SpaceRef Editor
May 29, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below. Day
174 on orbit for Expedition 4.

Countdown of STS-111/Endeavour (UF-2) continues for tomorrow
evening’s planned liftoff.  Weather prediction is 60%
probability of No Go due to the typical late afternoon chance of
thunderstorms each day, with further deterioration to 70% for Friday.
 If Friday is No Go, Endeavour will stand down for cryo
reservicing and then try again not earlier than Monday, 6/3.

MCC-H and MCC-M are in the final process of developing contingency
plans for a UF-2 launch slip from tomorrow evening all the way to a
nine-day delay.  Depending on the number of days of launch slip,
alternative workday timelines are being designed to replace the
scheduled joint Shuttle/ISS flight plan activities.  This
rescheduling process, which also requires adjustment of sleep shift
cycles, is affected by how quickly the decision to slip the launch
occurs, the magnitude of the launch delay, and any additional slips
which may occur after the initial launch slip.

Onboard ISS, preparations continue for the arrival of the new station
crew, and everything is getting "ship shape".  Storage
for UF-2 cargo has been prepared with great care (but it still will
represent a huge challenge).  EMU spacesuits and equipment are
all checked out and ready. Prepacking has gone well.

Also, the crew did a great job with yesterday’s pre-launch checkout
of the MSS (mobile service system)/Canadarm2.  Everything
checked out well, except for one minor anomaly.  [During the
JEU (joint electronics unit) diagnostics test for the Redundant
string, the SP (shoulder pitch) motor drive and windings test failed,
and failed again on repetition.  After driving the SP in both
coarse and vernier rates, the crew, on ground advice, repeated the
JEU diagnostics, and this time all the tests — including for the SP
 passed OK. It is believed that cycling the SSRMS brakes
corrected the situation.  The SSRMS/Robotics is considered ready
for UF-2.]

This morning, CDR Yuri Onufrienko began his day (2:10 am EDT)
with the conclusion of the acoustic environment survey started
yesterday in the two sleep compartments of the Service Module (SM).
 This morning, Yuri recorded the readings of both sound level
dosimeters for downlink during Russian AOS (acquisition of

For FE-1 Carl Walz, first daily duty was rebooting of PCSs (portable
computer systems), a daily routine task to ensure no locked-up
laptops, which is now regularly scheduled on the timeline.

Later, Onufrienko and Walz tagged up with Russian SUBA (onboard
equipment control system) specialists and performed additional cable
network outfitting on the SUBA, after the recent (5/21) installation
of another power control panel (PPS-26).  The task took several
hours. [They first unplugged three orthogonal angular rate sensors
(DUC) from the local analog commutator/switch unit (LKA) as well as
the respective PPS-26 connection to the BITS2-12 onboard telemetry
system, reconfigured the cabling, and then reconnected all
subsystems. SUBA is used to control, monitor and diagnose SM onboard
systems status by using sensor output signals and functional outputs
from the SM command radio link, onboard computer system (BVS) units,
SM control panels and onboard system relay outputs.  Its several
PPS consoles are used to switch electric power and to provide current
protection for the power supply circuits of separate instruments,
units, and devices.]

FE-2 Dan Bursch initiated EMU battery charging in the BSA
(battery stowage assembly), in which he had replaced chargers #3 and
#4 some time ago.  Thus, for the initial BSA charge, he powered
up only these two chargers, to give them a checkout. During a second
BSA charging round, all four chargers will be used.

Carl Walz performed regular maintenance on the CSA-CP (compound
specific analyzer/combustion products).  For its initial steps,
the checkout involved both CSA-CP units on the ER4 (EXPRESS Rack #4),
followed by a one-hour unattended data-logging period with only the
prime CSA-CP.

Daily routine servicing tasks were completed by Onufrienko (SOSH life
support maintenance) and Bursch (Inc. 4/8A payloads status checks,
IMS inventory "delta" file preparation).

All crewmembers performed their regular daily physical exercise on
TVIS (treadmill), CEVIS (cycle) and RED (resistive exerciser).
 For Yuri Onufrienko, the Russian flight surgeon has requested
some video of the CDR running on the TVIS with bungees as loading
devices.  [Dan Bursch was asked to shoot the videotape,
showing how Yuri gets onto the treadmill, straps himself into the
bungee harness (which presses him against the running surface), and
runs during the exercise, seen from various aspects of interest to
the flight surgeon.]

POC (Payload Operations Center) plans to return the BTR
(biotechnology refrigerator), currently in ER4, on STS-111.  The
contents of the arriving Stelsys CRIM (commercial refrigeration
incubator module) experiment, i.e., quad cells with human liver
preparation, will be power-transferred from the Shuttle middeck and
stored in the ARCTIC refrigerator/freezer.  [One of the
specialized functions of the liver is to break down drugs or toxins
into less harmful and more water-soluble substances that can be
excreted from the body. The StelSys bioreactor experiment will test
this function of human liver cells in microgravity vs. the function
of duplicate cells on Earth.]

Dan Bursch continued his "farming" work on the BPS
(biomass production system) test facility, today performing another
gas calibration and sampling on PGC4 (plant growth chamber #4).
 A total of five samples were taken,- four gas samples from PGC4
and one cabin air sample for calibration).

The originally scheduled correction of the BINS strapdown guidance
system using PUMA star fixes was cancelled due to the GPS system’s
recent return to full functionality for updating the BINS.

The crew had a tagup with the JSC Astronaut Office via

A final definitive answer to yesterday’s crew puzzler, why the Lab
consistently has a slightly higher CO2 concentration than the Russian
SM module, was uplinked to the crew: the Lab’s CDRA (carbon dioxide
removal assembly) is not running, and the SM’s Vozdukh is currently
the only CO2 remover. The further removed from Vozdukh, the higher
the CO2 will be (note: all within limits). If CDRA were running and
Vozdukh not, the US segment would show lower CO2 concentrations.

When the crew, on 5/24, replaced the active filtering element of the
Russian Potok-150MK (150 micron) microbial filtration system, they
attempted to jumper around a blown fuse, not having a replacement. A
power circuit card began smoking and burned through.  The power
unit needs to be replaced, possibly with a new unit brought up by
Progress 8P, and the damaged element returned to Earth, possibly on

Instead of the originally considered swap-out of the roles of GNC1
(guidance, navigation & control computer #1) and GNC2 for the
UF-2 docked stage, the plan now is to have a software PPL
(pre-programmed load) on board which would inhibit a handover to the
RS in the event of a power failure. As backup, Moscow will have a
second inhibit command ready for uplink. [Background: As reported
earlier (5/17), the GNC1, currently Primary, and the Primary C&C
MDM3 (command and control computer #3) are on the same power feed,
while GNC2 is on a separate channel.  Should the power fail,
this arrangement would deactivate both of the Primaries, and the
backup C&C would come up before GNC2, be unable to
"see" a running GNC, and immediately invoke an attitude
control handover to the Russian segment (RS).  If this occurs
after Shuttle docking, the SM motion control system would begin a
force-fight with the Orbiter’s attitude control systems, endangering

Yuri Onufrienko’s observational ocean target zones for the
Diatomeya experiment today were the Indian Ocean, the Atlantic Ocean
(SE coast of South America with La Plata Bay, Parana River drainage
and the waters above undersea canyons), and the Pacific Ocean (W of
Chile, areas of underwater volcanic activity).

Today’s U.S. CEO (crew earth observation) targets were
Industrialized SE Africa (a cold front moving in from the
south was preventing seaward movement of smoke and smog from Pretoria
and Johannesburg. Crew was to look right as they passed N of the
urban centers and record any aerosols lying in the Orange-Vaal
valley, as well as in local pockets)
, Patagonian Glaciers
(following the passage of a strong weather front, the crew should
have had an uncommonly good view of the W flank of the Andes, the
coastal fjords and islands. Of interest: photographs of the islands
and the heads of the fjords, which are usually cloud covered. Oblique
context views  were also needed of the South Patagonian ice
, Parana River (beginning at the right-angle bend
in the river near the city of Rosario, this ascending pass followed
the Parana to its headwaters. Looking right of track, the crew was to
take a mapping strip [10% overlap of frames] from Rosario to the
reservoirs in the coastal highlands)
, and Lake Poopo
(as ISS crossed the Altiplano near Lake Titicaca, of interest:
looking right for a view that embraces Lake Poopo and the salars
Uyuni and Coipasa. Sea surface temperatures are now slightly elevated
off the northwest coast of South America, although atmospheric
pressures and wind circulation patterns that would be confirming for
an El Nino event are not yet established).
CEO images can be viewed at the website

Update on yesterday’s conjunction report (Delta 2
rocket body, object #22658): TCA (time of closest approach: 5/28,
11:12 pm EDT. Miss distance (radial): 4.5 km. Never became a real
threat due to the large radial miss distance.

Conjunction (Unknown object, # 88838):  TCA: 5/29 at 2:10
pm EDT.  Radial miss distance: 10.0 km. Cross-sectional area:
0.004 sq.m. Plan of Action: Due to the late notification, maneuver
planning could not be performed.  This object is very small and
difficult to track.  US Spacecom has very few observations on
this object and very low confidence in their solution.

U.S. and Russian Segment Status (as of 1:46 pm EDT):

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode).
Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump
failed).  U.S. CDRA CO2 scrubber is Off.  BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify

SM Working Compartment:  Pressure (mmHg) — 759,
temperature (deg C) — 28.3, ppO2 (mmHg) — 148.8, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 762,
temperature (deg C) — 21.2.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 21.7.
Node:  Pressure (mmHg) — 750.89, temperature (deg C) —
23.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 752.49, temperature (deg C)
— 25.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 752.39,
temperature (deg C) — 22.4; shell heater temp (deg C) — 22.4, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 23.5
PMA-2:  Shell heater temp (deg C) — 11.4

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Autotrack mode.
SM batteries: Battery #2 is offline. Battery #7 is in Cycle mode. All
other batteries (6) are in “partial charge” mode.
FGB: Battery #3 is offline (ROM mode). All other batteries (5) are in
“partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating

SSRMS/Canadarm2 at Progress viewing position, with Keep Alive
power on both strings.
RWS (robotics workstations) are Off.
SSRMS Prime string Wrist Roll (WR) joint to be replaced on UF-2.

ISS Orbit (as of this morning, 6:47 am EDT [=
Mean altitude — 387.4 km
Apogee — 392.0 km
Perigee — 382.9 km
Period — 92.3 min.
Inclination (to Equator) — 51.63 deg
Eccentricity — 0.0006726
Orbits per 24-hr. day — 15.60
Altitude decrease — 150 m (mean) in last 24 hours
Solar Beta Angle — 71.4 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 20110
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: ~180 deg, pitch: -5.6 deg., roll: 0

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.